A titration was carried out to determine the concentration of a hydrogen peroxide solution. 5cm3 of the hydrogen peroxide solution was drawn off and diluted with water to 250cm3 solution in a volumetric flask. 25cm3 of this solution was pipetted into a conical flask and acidified. It was neutralised by 22.10cm3 of a potassium manganate(VII) solution of strength 0.0170 moldm–3. Calculate the concentration of the original hydrogen peroxide solution (8 marks)


1
Expert's answer
2021-09-11T00:04:39-0400

5/ 34.0147= 0.147mol

250/1000 = 0.25

= 0.147/0.25 = 0.58 mol/L